OpenStudy (anonymous):

1. How are the functions y = x and y = x – 3 related? How are their graphs related? (1 point)

OpenStudy (anonymous):

Each output for y = x – 3 is 3 more than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. Each output for y = x – 3 is 3 more than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated up 3 units. Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated up 3 units

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Teddyiswatshecallsme